rectangle with a length of 5 inches and width of 3 inches. if, the actual length is 50 feet, what is the scale factor?
Answer by macston(5194)

Scale factor=Actual measurement/scaled measurement=(50feet/5in)(12in/ft)=600/5=120/1
In ratio form, the scale factor is 120:1.